July 16, 2186 a solar eclipse lasting seven minutes and 29 seconds will be observed over the mid Atlantic ocean during the eclipse in observer may need a flashlight to see if the flashlight draws a current of .22 A how much charge will pass to the lightbulb during the eclipse
1
Expert's answer
2020-05-08T15:27:52-0400

The current strength shows what charge passes through the cross section of the conductor per unit time.

"I=\frac{q}{t}"

then write

"q=I \cdot t=0.22 \cdot(7 \cdot60+29)=0.22 \cdot 449=98.78 C"